Consider the reaction: 3HNO3(aq) + K3PO4(aq)  H3PO4 + 3KNO3(aq) What is the concentration of an HNO3 solution if 25.0 mL reacts with 38.5 mL of 0.145 M K3PO4?




1
Expert's answer
2021-11-22T14:57:03-0500

Solution:

Balanced chemical equation:

3HNO3(aq) + K3PO4(aq) → H3PO4 + 3KNO3(aq)

According to stoichiometry:

3 mol of HNO3 reacts with 1 mol of K3PO4

Therefore,

Moles of HNO3 = 3 × Moles of K3PO4


Moles = Molarity × Volume

Therefore,

Molarity of HNO3 × Volume of HNO3 = 3 × Molarity of K3PO4 × Volume of K3PO4

Molarity of HNO3 × (25.0 mL) = 3 × (0.145 M) × (38.5 mL)

Molarity of HNO3 = (3 × 0.145 M × 38.5 mL) / (25.0 mL) = 0.6699 M = 0.670 M

Molarity of HNO3 = 0.670 M


Answer: The molar concentration of an HNO3 solution is 0.670 M